OpenStudy (anonymous):

Multiply: (-5x + 4)(4x2 + x)

OpenStudy (anonymous):

(-5x + 4)(4x^2 + x)=(-5x + 4)(4x + 1)x= (-20x^2 - 5x + 16x +4)x= (-20x^2 + 11x +4)x=-20x^3 + 11x^2 +4x